OpenStudy (anonymous):

30. Find the vector v with the given magnitude and the same direction as u. which v=3 and u=(4,-4)

OpenStudy (anonymous):

So we want a vector, v, that looks like this:  where the y component is the negative x component

OpenStudy (anonymous):

if the magnitude is 3, then \[v_x^2 + v_y^2 = 3^2\] we also know that \[v_x =-v_y\]

OpenStudy (anonymous):

so: \[2v_x^2 = 9\]

OpenStudy (anonymous):

therefore: \[v_x = \sqrt{9/2}\]

OpenStudy (anonymous):

and: \[v_y = -\sqrt{9/2}\]